from the equation with rational coefficients whose roots are -1,3+2i?
Answer by dkppathak(439)

roots are -1,3+2i
sum of roots =-1+3+2i=2+2i
product of roots =-1x(2+2i)=-2-2i
equations
x^2+(sum of roots)+ product of roots
x^2+(2+2i)x-(2+2i)^2
x^2+(2+2i)x-8i
x^2+6ix+2x
answer
x^2+(2+6i)x